# VoxCity: A Seamless Framework for Open Geospatial Data Integration, Grid-Based Semantic 3D City Model Generation, and Urban Environment Simulation

### v1

@@ 1 Introduction
Meanwhile, increasing global geospatial data relevant to 3D city models has been openly released in the last few years, largely due to rapid advancements in deep learning techniques [22, 23, 24, 25]. For instance, Sirko et al. [26] released building height data with a 4 m spatial resolution for the Global South; Tolan et al. [23] published global canopy height data at a 1 m resolution; Zanaga et al. [24] released land cover data with worldwide coverage at a 10 m resolution; and Hawker et al. [25] created global terrain elevation data with a 1 Arc-second (30.9 m at the equator) resolution, based on Copernicus DEM, canopy height and building footprint data. We can use these open geospatial datasets to create 3D city models with semantic information regarding buildings, trees, land cover, and terrain. However, while these datasets can reduce the efforts required to collect necessary information and increase the coverage, they still require intensive manual efforts to integrate them and reconstruct 3D city models of sufficient quality for advanced uses such as simulations. Each dataset has different data types, such as raster and vector, and varying spatial resolutions, making the data integration process complex. To scale applications of urban environment simulation globally, it is desirable to automate these manual efforts for data integration. Moreover, even with simulation-ready models, working in simulation software remains time-consuming, particularly for tasks such as data import, boundary condition settings, and result data export.
@@
To bridge the gap between 3D model generation and urban environment simulation, a meshing process fundamentally needs to be incorporated. Many urban environment simulation methods include the process where 3D city models are split into meshes with certain scales. For example, Computational Fluid Dynamics (CFD) for wind environment simulation requires volume meshes for fluid (air) domains and surface meshes for solid objects, such as buildings and ground [29]. Ray-tracing for solar irradiance simulation requires surface meshes of solid objects, where solar radiation is reflected or absorbed [30]. There are two principal methods for the meshing process: structured meshing and unstructured meshing. Structured meshing is a grid-based approach to create cuboid volume meshes and rectangular surface meshes [31, 32]. Unstructured meshing does not rely on a grid and generates polyhedral volume meshes and triangular surface meshes, allowing for more flexibility in shape and proportion [33, 29]. Numerous software packages employ structured meshing due to its lower computational load for both calculation and memory, as well as its stability in numerical simulations. Unstructured meshing has the advantage of incorporating curved surfaces; however, it can generate low-quality meshes with overly sharp or blunt vertex angles, leading to divergence or low accuracy in numerical simulations. For city-scale applications, structured meshing should be more desirable than unstructured meshing. More specifically, the voxel-based approach, which applies the same mesh size for the entire domain in structured meshing[34, 35], offers advantages in simplicity, as a single 3D array and a specified voxel size can represent the entire mesh. Moreover, voxels have better compatibility with raster formats that are commonly used in geospatial data, such as land cover, tree canopy height, and terrain elevation.
@@ 2.1 Urban Environment Simulation
Climate change [36] has intensified heat-related challenges like frequent heat waves [37, 38] and urban heat island effects [39], spurring increased assessment of outdoor heat stress and thermal comfort through CFD [40, 2], radiation transfer models [1, 41, 42], and heat conduction and energy balance analyses [43]. The objectives of microclimate simulations have encompassed not only heat-related health issues but also other diverse aspects such as energy consumption for heating, ventilation, air conditioning (HVAC), and lighting [6, 7, 44, 45, 46], air quality [27, 8], wind comfort [47, 48], wind pressure on building surfaces [49, 50], and sunlight and ultraviolet exposure focusing on human skin health [51]. Additionally, numerous studies have reported impacts of buildings [52, 53, 54, 55], trees [56, 57, 58, 59, 60], low vegetation [61, 62, 63], water bodies [64, 65, 66, 67, 68], and terrain morphology [69] on the microclimate, indicating the importance of including these urban elements in 3D city models for simulations.
Meanwhile, many studies have conducted urban environment simulations to assess the visual comfort of streetscapes [4, 5] and window views [70]. Specifically, sky view factor (SVF) or sky view index (SVI) [71] and green view index (GVI) [72] have mainly been evaluated as quantitative indicators using 3D city models with buildings, trees, and low vegetation land covers. Numerous studies have employed street view imagery and computer vision techniques instead of simulations with 3D city models to evaluate view indices [73, 74, 75, 76]. This enables view analysis without preparing detailed 3D city models; however, such analysis cover only viewpoints along street networks, and furthermore, not all streets have sufficient imagery. These view indicators have been assessed focusing not only on visual comfort but also on walkability [77, 74] and bikeability [78] of streets, the well-being of city dwellers [79], and the preference of residents and real-estate prices [80, 81]. The visibility of water features, including the ocean, lakes, and rivers, has also been discussed in the same context [82].
Our literature review revealed that buildings, trees, land cover, water bodies, and terrain elevation have been included as essential information in 3D city models across different simulation objectives.
@@ 2.2 Creation of 3D city models
Meanwhile, an increasing number of studies have employed open 3D city models that have been individually released by countries or cities [86, 87, 88, 89, 90]. However, in most cases, we still need to create 3D city models ourselves because cities with such available models are limited [19]. Furthermore, the data, even when openly available, is not always directly compatible with urban environment simulations. For instance, while 3D city models fundamentally include building data, only a portion of them have other semantic information, including trees and land cover [20]. Many software solutions for urban environment simulation require their own input file formats for 3D city models, such as INX for ENVI-met [^1] https://envi-met.com/ and MSH for ANSYS [^2] https://www.ansys.com/. Additionally, some simulations require high completeness in geometry for their meshing processes. Software for CFD simulations tends to require precise watertight geometries for solid objects and does not accept polygons with gaps and intersections [3, 21]. Therefore, even for studies targeting areas with open 3D city models, many of them need to add missing information from other data sources, convert file formats, or modify to meet quality standards.
@@ 2.3 Related work
There are several open tools to generate 3D city models. For instance, 3dfier [27] is an open-source software to generate 3D building models from point cloud data and building footprints. BlenderGIS [^3] https://blender-addons.org/blendergis-addon/ is a plug-in tool for Blender [^4] https://www.blender.org/, a free and open-source 3D modeling software, which generates 3D city models by combining building data from OpenStreetMap [^5] https://www.openstreetmap.org and terrain elevation from Shuttle Radar Topography Mission (SRTM) [91]. However, we were unable to find any open-source tools that could generate 3D city models containing all essential elements and attributes, including buildings, vegetation, water bodies, and terrain.
@@
Several previous studies suggest a potential solution to the demand for a semantic 3D city model generator. Ding et al. [92] and Li and Wang [42] integrated multiple geospatial data to create 3D city models with semantic information and conducted microclimate simulations using the generated models. Specifically, Ding et al. [92] combined data on building height, land cover, and terrain elevation, while Li and Wang [42] combined building height and tree canopy height. Although Ding et al. [92] and Li and Wang [42] created land cover data using aerial imagery and building data using aerial LiDAR data respectively, they directly use openly available data sources for other data types. If these methods can be extended to encompass all essential data types—building height, land cover, canopy height, and terrain elevation—and if all these data can be collected from openly and globally available data sources, such a tool would be highly beneficial for many researchers and engineers working on urban environment simulations.
@@
Existing open datasets vary considerably in their spatial coverage, accuracy, and resolutions. Despite the importance of these factors for application in 3D city modeling and urban environment simulation, we were unable to find any studies offering a systematic comparison for such datasets. This gap underscores the importance of cataloging available open geospatial data, evaluating their quality, and their applicability in generating semantic 3D city models. However, no studies have reviewed open geospatial data from this perspective.
@@
Meanwhile, increasing studies have proposed open-source tools as Python packages [27, 93, 94, 95, 96, 97]. For example, Yap et al. [93] introduced the Python package ‘Urbanity’ to automate the construction of feature-rich urban networks at any geographical scale and any location. Gholami et al. [94] developed a Python-based approach to assess microscale human thermal stress in urban environments. Python is one of the most dominant programming languages for data science, especially in machine learning. Many Python users openly share their scripts and packages through platforms such as GitHub [^6] https://github.com/ and PyPI [^7] https://pypi.org/, facilitating easier and more scalable distribution of developed tools.
@@
Therefore, we propose a tool that offers three key contributions: reducing the manual effort required to prepare 3D city models for urban environment simulations, and enabling broader applications of these simulations in environmental research and urban-architectural design projects. To achieve these contributions, we present three main developments:
@@ 3 Review of open geospatial data
In this section, we review open geospatial data related to building height, tree canopy height, land cover, and terrain elevation, comparing spatial coverage, resolution, platform, and file format. Sections 3.1 to 3.4, discuss each data type in detail. Based on the review, we present a comprehensive data catalog in Section 3.5 to guide readers in selecting suitable data sources for specific target areas and research purposes.
@@ 3.1 Building height
We first screened open building height datasets using two criteria: (1) global or multi-national coverage and (2) horizontal resolution finer than 10 m. The 10-meter threshold ensures sufficient granularity to represent urban morphology, including small-scale buildings and street canyons. During the screening process, we excluded datasets that cover only one country or city [98, 99, 100, 101] or have coarser than 10 m resolution [102, 103, 104]. As a result, we selected OpenStreetMap [^8] https://www.openstreetmap.org (OSM), Overture [^9] https://overturemaps.org, EUBUCCO v0.1101010https://eubucco.com/ [105], UT-GLOBUS [22], Open Buildings 2.5D Temporal dataset (OB2.5DT) [26], and Microsoft Building Footprints (MSBF) [106]. Table 1 summarizes their characteristics. None of these sources provides complete, worldwide coverage, underscoring the importance of selecting a dataset that best suits the target city.
Table 1: Comparison of building height data sources.
@@
OpenStreetMap Worldwide (24% completeness in city centers [107]) - / Not provided API / JSON (vector) Volunteered / updated continuously
Overture Worldwide - / Not provided API / JSON (vector) OpenStreetMap, Esri Community Maps Program, Google Open Buildings, etc. / updated continuously
@@
Overall, OSM, Overture, and EUBUCCO demonstrate sufficient quality for many applications, whereas MSBF, UT-GLOBUS, and OB2.5DT are less suited as sole sources for 3D city model generation. They can, however, serve as complementary data for missing building height in EUBUCCO, OSM, and Overture. In Figure 2, the panels labeled “OSM + [dataset]” and “Overture + [dataset]” illustrate the integration of OSM or Overture footprints with either MSBF, UT-GLOBUS, or OB2.5DT. Where OSM or Overture footprints lack a height value, the missing attribute is retrieved from the intersecting footprints in a complementary data source. If a building intersects multiple footprints, the final height is determined by a weighted average based on intersection area. The resulting combined data exhibit more complete height coverage than OSM or Overture alone.
@@ 3.2 Land cover
We filtered open land cover datasets using the following criteria: (1) worldwide coverage, (2) land cover classes suitable for general urban environment simulations, and (3) horizontal resolution finer than 10 m. We excluded datasets that cover only a single city or country [109, 110], are restricted to specific land cover types (e.g., ice, crop, or forest) [111, 112, 113, 114], or have resolution coarser than 10 m [115, 116, 117].
As a result, we selected ESA World cover (ESA) [24], ESRI 10m Annual Land Cover (2017-2023) (Esri) [118], Dynamic World V1 (DW) [119], OpenStreetMap [^11] https://www.openstreetmap.org (OSM), OpenEarthMap Japan (OEMJ) [120], and UrbanWatch (UW) [121] for our further review. Although OEMJ and UW each cover only one nation, we included them because of their exceptionally high (1 m) resolution. Table 2 compares key metrics for these sources.
@@
Dynamic World V1 [119] Water, Trees, Grass, Flooded vegetation, Crops, Shrub and scrub, Built, Bare, Snow and ice Worldwide 10 m / 73.8% Google Earth Engine, Zenodo / GeoTIFF (Raster) Prediction from satellite imagery / updated continuously
OpenStreetMap Bare rock, Rock, Sand, Desert, Grass, Park, Industrial, Construction, Railway, Parking, Highway, Wood, Forest, Tree, Water, Waterway, Bay, Ocean, Farmland, Building, etc. Worldwide - / Not provided API / JSON (vector) Volunteered / updated continuously
@@
Figure 3 shows example land cover maps from the reviewed datasets for Osaka, Los Angeles, and São Paulo. To allow direct comparisons, we harmonized the class definitions across all datasets using the conversion rules presented in Table 3. Overall, OEMJ (for Osaka) and UW (for Los Angeles) most closely match the actual land cover when compared to aerial imagery. OSM also aligns well but tends to miss smaller patches of vegetation (particularly trees). ESA captures major water bodies and broad vegetation patterns but is less detailed than OEMJ, UW, or OSM. In OSM, some areas (e.g., São Paulo) lack detail in spatial variation, indicating lower completeness than ESA. ESRI and DW indicate low completeness, classifying most pixels as “Developed space” despite these areas actually containing diverse classes. We attribute the relatively low performance of ESRI and DW to their primary focus not being on urban areas, resulting in training and validation data that lacks detailed annotations for urban regions.

### v3

@@ Abstract
This is the Accepted Manuscript version of an article published by Elsevier in the journal Computers, Environment and Urban Systems in 2026, which is available at: https://doi.org/10.1016/j.compenvurbsys.2025.102366 Cite as: Fujiwara K, Tsurumi R, Kiyono T, Fan Z, Liang X, Lei B, Yap W, Ito K, Biljecki F (2026): VoxCity: A seamless framework for open geospatial data integration, grid-based semantic 3D city model generation, and urban environment simulation. Computers, Environment and Urban Systems, 123: 102366.
© 2026, Elsevier. Licensed under the Creative Commons Attribution-NonCommercial-NoDerivatives 4.0 International (http://creativecommons.org/licenses/by-nc-nd/4.0/)
@@ 1 Introduction
Meanwhile, an increasing amount of global geospatial data relevant to 3D city models has been openly released in the last few years, largely due to rapid advancements in deep learning techniques [22, 23, 24, 25]. For instance, Sirko et al. [26] released building height data with a 4 m spatial resolution for Africa, South Asia, Southeast Asia, Latin America and the Caribbean; Tolan et al. [23] published global canopy height data at a 1 m resolution; Zanaga et al. [24] released land cover data with worldwide coverage at a 10 m resolution; and Hawker et al. [25] created global terrain elevation data with a 1 Arc-second (30.9 m at the equator) resolution, based on Copernicus DEM, canopy height and building footprint data. We can use these open geospatial datasets to create 3D city models with semantic information regarding buildings, trees, land cover, and terrain. However, while these datasets can reduce the efforts required to collect necessary information and increase the coverage, they still require intensive manual efforts to integrate them and reconstruct 3D city models of sufficient quality for advanced uses such as simulations. Each dataset has different data types, such as raster and vector, and varying spatial resolutions, making the data integration process complex. To scale applications of urban environment simulation globally, it is desirable to automate these manual efforts for data integration. Moreover, even with simulation-ready models, working in simulation software remains time-consuming, particularly for tasks such as data import, boundary condition settings, and result data export.
@@
To bridge the gap between 3D model generation and urban environment simulation, a meshing process fundamentally needs to be incorporated. Many urban environment simulation methods include the process where 3D city models are split into meshes with certain scales. For example, Computational Fluid Dynamics (CFD) for wind environment simulation requires volume meshes for fluid (air) domains and surface meshes for solid objects, such as buildings and ground [29]. Ray-tracing for solar irradiance simulation requires surface meshes of solid objects, where solar radiation is reflected or absorbed [30]. There are two principal methods for the meshing process: structured meshing and unstructured meshing. Structured meshing is a grid-based approach to create cuboid volume meshes and rectangular surface meshes [31, 32]. Unstructured meshing does not rely on a grid and generates polyhedral volume meshes and triangular surface meshes, allowing for more flexibility in shape and proportion [33, 29]. Numerous software packages employ structured meshing due to its lower computational load for both calculation and memory, as well as its stability in numerical simulations. Unstructured meshing has the advantage of incorporating curved surfaces; however, it can generate low-quality meshes with overly sharp or blunt vertex angles, leading to divergence or low accuracy in numerical simulations. For city-scale applications, structured meshing should be more desirable than unstructured meshing. More specifically, the voxel-based approach, which applies the same mesh size for the entire domain in structured meshing [34, 35, 36], offers advantages in simplicity, as a single 3D array and a specified voxel size can represent the entire mesh. Moreover, voxels have better compatibility with raster formats that are commonly used in geospatial data, such as land cover, tree canopy height, and terrain elevation.
@@ 2.1 Urban Environment Simulation
Climate change [37] has intensified heat-related challenges like frequent heat waves [38, 39] and urban heat island effects [40], spurring increased assessment of outdoor heat stress and thermal comfort through CFD [41, 2], radiation transfer models [1, 42, 43], and heat conduction and energy balance analyses [44]. The objectives of microclimate simulations have encompassed not only heat-related health issues but also other diverse aspects such as energy consumption for heating, ventilation, air conditioning (HVAC), and lighting [6, 7, 45, 46], air quality [47, 48], wind comfort [49, 50], wind pressure on building surfaces [51, 52], and sunlight and ultraviolet exposure focusing on human skin health [53]. Additionally, numerous studies have reported impacts of buildings [54, 55, 56, 57], trees [58, 59, 60, 61, 62], low vegetation [63, 64, 65], water bodies [66, 67, 68], and terrain morphology [69] on the microclimate, indicating the importance of including these urban elements in 3D city models for simulations.
Meanwhile, many studies have conducted urban environment simulations to assess the visual comfort of streetscapes [5] and window views [70, 71]. Specifically, sky view factor (SVF) or sky view index (SVI) [72] and green view index (GVI) [73, 71, 74] have mainly been evaluated as quantitative indicators using 3D city models with buildings, trees, and low vegetation land covers. Numerous studies have employed street view imagery and computer vision techniques instead of simulations with 3D city models to evaluate view indices [75, 76, 77, 78, 79]. This enables view analyses without preparing detailed 3D city models; however, such analyses cover only viewpoints along street networks, and furthermore, not all streets have sufficient imagery. These view indicators have been assessed focusing not only on visual comfort but also on walkability [80, 76] and bikeability [81] of streets, the well-being of city dwellers [82], and the preferences of residents and real-estate prices [83, 84]. The visibility of water features, including the ocean, lakes, and rivers, has also been discussed in the same context [85].
Table 7 in Appendix illustrates the relationships between simulation categories and the urban elements incorporated into 3D city models used in previous studies. Our literature review revealed that buildings, trees, land cover, water bodies, and terrain elevation have been included as essential information in 3D city models across different simulation objectives.
@@ 2.2 Creation of 3D city models
Meanwhile, an increasing number of studies have employed open 3D city models that have been individually released by countries or cities [89, 90, 91, 92, 93]. However, in most cases, we still need to create 3D city models ourselves because cities with such available models are limited [19]. Furthermore, the data, even when openly available, is not always directly compatible with urban environment simulations. For instance, while 3D city models fundamentally include building data, only a portion of them have other semantic information, including trees and land cover [20]. Many software solutions for urban environment simulation require their own input file formats for 3D city models, such as the ENVI-met [^1] https://envi-met.com/ geometry file (INX) and the ANSYS [^2] https://www.ansys.com/ mesh file (MSH). Additionally, some simulations require geometrically complete models for their meshing processes. Software for CFD simulations tends to require precise watertight geometries for solid objects and does not accept polygons with gaps and intersections [3, 21]. Therefore, even for studies targeting areas with open 3D city models, many of them need to add missing information from other data sources, convert file formats, or modify them to meet quality standards.
@@ 2.3 Related work
There are several open-source tools to generate 3D city models. For instance, 3dfier [27] is an open-source software to generate 3D building models from point cloud data and building footprints. BlenderGIS [^3] https://blender-addons.org/blendergis-addon/ is a plug-in tool for Blender [^4] https://www.blender.org/, a free and open-source 3D modeling software, which generates 3D city models by combining building data from OpenStreetMap [^5] https://www.openstreetmap.org and terrain elevation from Shuttle Radar Topography Mission (SRTM) [94]. However, we were unable to find any open-source tools that could generate 3D city models containing all essential elements and attributes, including buildings, vegetation, water bodies, and terrain.
@@
Several previous studies suggest a potential solution to the demand for a semantic 3D city model generator. Ding et al. [95] and Li and Wang [43] integrated multiple geospatial data to create 3D city models with semantic information and conducted microclimate simulations using the generated models. Specifically, Ding et al. [95] combined data on building height, land cover, and terrain elevation, while Li and Wang [43] combined building height and tree canopy height. Although Ding et al. [95] and Li and Wang [43] created land cover data using aerial imagery and building data using aerial LiDAR data respectively, they directly used openly available data sources for other data types. If these methods can be extended to encompass all essential data types—building height, land cover, canopy height, and terrain elevation—and if all these data can be collected from openly and globally available data sources, such a tool would be highly beneficial for many researchers and engineers working on urban environment simulations.
@@
Existing open datasets vary considerably in their spatial coverage, accuracy, and resolutions. Despite the importance of these factors for application in 3D city modeling and urban environment simulation, we were unable to find any studies offering a systematic comparison for such datasets. This gap underscores the importance of cataloging available open geospatial data, and evaluating their quality and applicability in generating semantic 3D city models. However, no studies have reviewed open geospatial data from this perspective.
@@
Meanwhile, an increasing number of studies have proposed open-source tools as Python packages [27, 96, 97, 98, 99, 100]. For example, Yap et al. [96] introduced the Python package ‘Urbanity’ to automate the construction of feature-rich urban networks at any geographical scale and any location. Gholami et al. [97] developed a Python-based approach to assess microscale human thermal stress in urban environments. Python is one of the most dominant programming languages for data science, especially in machine learning. Many Python users openly share their scripts and packages through platforms such as GitHub [^6] https://github.com/ and PyPI [^7] https://pypi.org/, facilitating easier and more scalable distribution of developed tools.
@@
Therefore, we propose a tool that offers two key contributions: reducing the manual effort required to prepare 3D city models for urban environment simulations, and enabling broader applications of these simulations in environmental research and urban-architectural design projects. To achieve these contributions, we present three main developments:
@@ 3 Review of open geospatial data
In this section, we review open geospatial data related to building height, tree canopy height, land cover, and terrain elevation, comparing spatial coverage, resolution, platform, and file format. Sections 3.1 to 3.4 discuss each data type in detail. To select data sources to review, we set two criteria: (1) global or multinational coverage and (2) horizontal resolution finer than 10 m. We set this 10 m threshold, following previous studies that employ 10 m as the maximum mesh size for wind and microclimate simulation [101, 102, 103]. Additionally, our data review has not incorporated existing 3D city models such as Project PLATEAU for Japan [^8] https://www.mlit.go.jp/plateau/open-data/ because this work specifically focuses on contributing to cities without such available models. This limitation is discussed in Section 5.1. Based on this review, we present a comprehensive data catalog in Section 3.5 to guide readers in selecting suitable data sources for specific target areas and research purposes.
@@ 3.1 Building height
Following the aforementioned criteria for spatial coverage and resolution, we selected OpenStreetMap [^9] https://www.openstreetmap.org (OSM), Overture [^10] https://overturemaps.org, EUBUCCO v0.1111111https://eubucco.com/ [104], UT-GLOBUS [22], Open Buildings 2.5D Temporal dataset (OB2.5DT) [26], and Microsoft Building Footprints (MSBF) [105] for review. Table 1 summarizes their characteristics. None of these sources provides complete, worldwide coverage, underscoring the importance of selecting a dataset that best suits the target city.
Table 1: Comparison of building height data sources. RMSE and MAE denote Root Mean Squared Error and Mean Absolute Error, respectively.
@@
OpenStreetMap Worldwide (24% completeness in city centers [106]) - / Not provided API / JSON (vector) Volunteered / updated at irregular intervals
Overture Worldwide - / Not provided API / JSON (vector) OpenStreetMap, Esri Community Maps Program, Google Open Buildings, etc. / updated at irregular intervals
@@
Overall, OSM, Overture, and EUBUCCO demonstrate sufficient quality for many applications, whereas MSBF, UT-GLOBUS, and OB2.5DT are less suited as sole sources for 3D city model generation. They can, however, serve as complementary data for missing building heights in EUBUCCO, OSM, and Overture. In Figure 2, the panels labeled “OSM + [dataset]” and “Overture + [dataset]” illustrate the integration of OSM or Overture footprints with either MSBF, UT-GLOBUS, or OB2.5DT. Where OSM or Overture footprints lack a height value, the missing attribute is retrieved from the intersecting footprints in a complementary data source. If a building intersects multiple footprints, the final height is determined by a weighted average based on intersection area. The resulting combined data exhibit more complete height coverage than OSM or Overture alone.
@@ 3.2 Land cover
To select data sources for review, we set criteria requiring land cover classes suitable for general urban environment simulations, in addition to the aforementioned criteria for spatial coverage and resolution. Therefore, we excluded datasets that are restricted to specific land cover types (e.g., ice, crop, or forest) [108, 109, 110, 111].
As a result, we selected ESA World Cover 10 m 2021 V200 (ESA) [24], Esri 10m Annual Land Cover (2017-2023) (Esri) [112], Dynamic World V1 (DW) [113], OpenStreetMap [^12] https://www.openstreetmap.org (OSM), OpenEarthMap Japan (OEMJ) [114], and UrbanWatch (UW) [115] for our further review. Although OEMJ and UW each cover only one nation, we included them because of their exceptionally high (1 m) resolution. Table 2 compares key metrics for these sources.
@@
Dynamic World V1 [113] Water, Trees, Grass, Flooded vegetation, Crops, Shrub and scrub, Built, Bare, Snow and ice Worldwide 10 m / 73.8% Google Earth Engine, Zenodo / GeoTIFF (Raster) Prediction from satellite imagery / updated daily
OpenStreetMap Bare rock, Rock, Sand, Desert, Grass, Park, Industrial, Construction, Railway, Parking, Highway, Wood, Forest, Tree, Water, Waterway, Bay, Ocean, Farmland, Building, etc. Worldwide - / Not provided API / JSON (vector) Volunteered / updated at irregular intervals
@@
Figure 3 shows example land cover maps from the reviewed datasets for Osaka, Los Angeles, and São Paulo. To allow direct comparisons, we harmonized the class definitions across all datasets using the conversion rules presented in Table 3. Overall, OEMJ (for Osaka) and UW (for Los Angeles) most closely match the actual land cover when compared to aerial imagery. OSM also aligns well but tends to miss smaller patches of vegetation (particularly trees). ESA captures major water bodies and broad vegetation patterns but is less detailed than OEMJ, UW, or OSM. In OSM, some areas (e.g., São Paulo) lack detail in spatial variation, indicating lower completeness than ESA. Esri and DW indicate low completeness, classifying most pixels as “Developed space” despite these areas actually containing diverse classes. We attribute the relatively low performance of Esri and DW to their primary focus not being on urban areas, resulting in training and validation data that lacks detailed annotations for urban regions. It is important to note that, while this is not applicable to the examples in Figure 3, OSM sometimes misses the water class for oceans because some ocean areas in OSM lack water object patches.